OpenStudy (anonymous):

tanx(1+cot^(2)x)=((1)/(cosxsinx))

OpenStudy (loser66):

prove?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

I am in hurry, I don't have time to walk you through, would you mind if I just give you the answer?

OpenStudy (anonymous):

i already know the answer., im just trying to prove it i got to 1/cotx(csc^(2)x)

OpenStudy (loser66):

cot = 1/tan--> cot^2 = 1/tan^2 apply tan( 1+ cot^2 ) = tan + cot^2 * tan = tan + 1/tan replace tan = sin/cos you have sin/cos + 1/(sin/cos) = sin/cos + cos/sin make them the same denominator \[\frac{sin}{cos}+\frac{cos}{sin}=\frac{sin^2 + cos^2}{cos*sin}=\frac{1}{cos * sin}\]

OpenStudy (loser66):

have to go. if you don't understand, ask others, please